OpenStudy (anonymous):

Please help me prove sec^2xsin^2x=sec^2x-1

OpenStudy (anonymous):

is it squared or just 2x?

OpenStudy (campbell_st):

well use the trig identity \[\sin^2(x) + \cos^2(x) = 1\] make the subject \[\sin^2(x)\] and then substitute into the left hand side of the equation. You also need to know \[\sec^2(x) = \frac{1}{\cos^2(x)}\] and that also may help

OpenStudy (schrodingers_cat):